Find the value of the finite population correction factor given the following N=3,000 and N=360

1
Expert's answer
2022-05-19T17:02:58-0400
"\sqrt{\dfrac{N-n}{N-1}}=\sqrt{\dfrac{3000-360}{3000-1}}\approx0.93824"